# Background resilient quantitative phase microscopy using entangled photons

### v1

@@ I Introduction
The use of entangled photons has also been extended to phase imaging. Phase imaging plays a pivotal role in revealing with enhanced clarity, the subtle variations in the optical path length of light passing through a transparent sample and has found widespread usage in biomedical science Park et al. (2018); Micó et al. (2019); Nguyen et al. (2022). Interferometry based quantum phase retrieval techniques includes the use of photonic N00N states Ono et al. (2013); Israel et al. (2014) and the interference between successive laser passes of spontaneous parametric down-conversion (SPDC) events Black et al. (2023). Phase imaging has also been demonstrated through holography with entangled photons Defienne et al. (2021b); Thekkadath et al. (2023). Non-interference methods include the use of transport of intensity equation between intensity images obtained from different image planes Lu et al. (2015); Ortolano et al. (2023) and through the simultaneous measurement of the position and momentum information of photons through a Fourier Ptychography approach Aidukas et al. (2019) and differential phase gradient approach Hodgson et al. (2023). These techniques offers various advantages such as sub-shot-noise sensitivity Ono et al. (2013); Israel et al. (2014); Ortolano et al. (2023); Lu et al. (2015), strong background-noise tolerance Aidukas et al. (2019); Defienne et al. (2021b) and ease of use, such as phase stability Thekkadath et al. (2023) or allowing scanning-free measurement compared to the classical counterpart Hodgson et al. (2023).
@@ II.1 Concept
The setup of our quantum phase microscope is depicted in Fig. 1. We utilize position-momentum entangled photon pairs, denoted as “signal” and “idler” photons, with orthogonal polarization generated through SPDC, to illuminate a phase target. After separating these two photons using a polarizing beamsplitter, the near-field (NF) of the target is imaged through a 4f imaging system by detecting the signal photons on a time-tagging camera, which records the detection time and location of each photon. Simultaneously, the far-field (FF) of the target, which contains momentum/angular information, is imaged by the idler photons onto a different region of the same camera. Knowing each photon pair is created at the same instance in time, they are identified through a time-correlation measurement using the detection time information from the camera. Because the photons are position-momentum entangled, we can infer the momentum of each signal photon detected at a NF position from the partnered idler photon, whose momentum is measured in the FF. Thus, 5-dimensional information, in time, position and momentum, is obtained for each photon.
@@
Note that background contributions has not been taken into account in eq. 1, see the Method section on the modifications required for background correction.
II.2 Quantitative phase measurement
@@ II.3 Background mitigation
Here, we simulate the effect of a dynamic background by introducing numerically generated photon detection events from a Gaussian shaped beam profile into the raw camera data, both in the NF and FF. The properties of this background, including its intensity, width, and position, were all randomized over time. Refer to the Methods section for more details on the background generation.
@@
However, even after coincidence detection, an uncorrelated background remains due to photons that accidentally arrive within the coincidence window despite not being generated in the same SPDC event. These are due to accidental coincidences between a SPDC photon and a background photon, two background photons, or two SPDC photons belonging to different pair events. In the second stage, the rate of accidental coincidences are measured by selecting a coincidence window that is outside of the true coincidence window, in this case, 50 ns before it. The contribution of background coincidences can then be subtracted from the the centroid shift to mitigate the effect of a dynamic background on the recovered phase profile. Refer to the Methods section for more details on the detection and subtraction of background light.
@@
II.4 Reduced far-field resolution
@@ III Discussion
IV Method
IV.1 Experimental Setup
A 20 mW, 405 nm continuous-wave (CW) laser, with a 1 mm beam diameter, is used to pump a 1 mm thick Type II periodically-poled potassium titanyl phosphate (ppKTP) crystal to generate, through the process of spontaneous parametric down-conversion (SPDC), orthogonally polarized photon pairs at 810 nm that are correlated in time and entangled in the position-momentum degrees of freedom. The photon pairs are directed through a 4f-imaging system to illuminate a phase target placed in the NF plane of the nonlinear crystal. The photons are then separated using a polarizing beamsplitter. The NF of the phase target is imaged onto a time-tagging camera (TPX3CAM Nomerotski (2019); ASI ) through the signal photons, while the FF of the target is projected onto a corner of the camera through a separate path using the idler photons.
To identify photon pairs, time correlation measurement is conducted with a coincidence window of 20 ns. Subsequently, a phase gradient measurement, as explained in the Concept section of the main text, is performed on each NF pixel to reconstruct the phase. Before conducting these measurements with the phase target, a one-time reference measurement without the phase target in place needs to be performed to determine $\phi_{o}(x,y)$.
IV.2 Background generation, detection and subtraction
Background photon events are numerically generated and inserted into the raw camera data. A random number generator is used to generate the detected position and time for each photon from a background light with Gaussian beam profile exhibiting Poissonian photon number statistics. This background is applied to both the NF and FF. Parameters for the intensity, width, and location of the Gaussian shaped background is randomized every 25 s to simulate a dynamic background. The background is chosen to be randomized every 25 s and not at a much faster rate so as not to create an overall background that is uniform in intensity after the 500 s data acquisition time. For the data in Fig. 4(b), the background intensity in each image plane is on average $\sim 0.5\times 10^{6}$ photons per second, and for Fig. 4(c) this is $\sim 1\times 10^{6}$ photons per second. In comparison, the number of detected SPDC photons in each image plane is $\sim 1.1\times 10^{6}$ photons per second. For better comparison, the same seed is used in the random number generation of the width and location for the background light at the two different intensity levels. An illustration on how the background is added to the NF and FF with time is shown in Fig. 6(a).
Figure 6: Background generation and detection (a) Shows the numerically added background light to the NF and FF whose parameters are randomized every 25 s. (b) A typical coincidence histogram taken showing the number of events detected as a function of the difference in arrival time between the photons detected on the two regions of the camera. The location and width of the coincidence window for identifying all photons detected in coincidence and only that of background coincidences are highlighted in orange. (c) Verifying the validity of the background identification approach illustrated in (b) through subtracting the obtained background image from the image containing all coincidences.
Figure 6(b) is a typical histogram showing the difference in arrival time between two photons detected in the two regions of the camera. The central peak is the result of time-correlated SPDC photon pairs being detected at the same time. On the other hand a constant background is also visible in the histogram, this is from the detection of coincidence between uncorrelated photons. It is not possible to determine which of the photon pairs detected are from background light, however, one can determine the resultant contribution of the background. To obtain the background contribution, the coincidence window is shifted away from the central peak. All photon pair events detected within this window will be from only uncorrelated photons. Since the background coincidence rate is uniform in time on the nanosecond timescale, this measurement is representative of the accidental coincidence rate in the central window. It can be verified that the resultant image formed by the uncorrelated photon pairs is indeed the background by subtracting it directly from the coincidence image to obtain a background free image, as shown in Fig. 6(c).
Due to the timing resolution of the camera, a 50 ns shift in the coincidence window is used in this experiment. For a detector with better timing resolution, which will result in a narrower central peak for SPDC, a smaller shift can be used. Note that this method of background identification is only valid when the background fluctuation is slower than the applied shift to the coincidence window.
Now with the position information on each detected photon pair from the background correlation measurement, the background contribution is subtracted from the centroid shift by modifying eq. 1 as follows:
$\mathscr{U}(x,y)$ $=\frac{1}{N(x,y)-N_{b}(x,y)}\left(\sum_{n=1}^{N(x,y)}u_{n}(x,y)-\sum_{m=1}^{N_{b}(x,y)}u^{b}_{m}(x,y)\right)$
$\mathscr{V}(x,y)$ $=\frac{1}{N(x,y)-N_{b}(x,y)}\left(\sum_{n=1}^{N(x,y)}v_{n}(x,y)-\sum_{m=1}^{N_{b}(x,y)}v^{b}_{m}(x,y)\right),$ (5)
where $N_{b}(x,y)$ is the number of background coincidences detected at position $(x,y)$ in the NF and $u^{b}(x,y)$ and $v^{b}(x,y)$ are the positions of the background photons detected in the FF, in coincidence with those detected at position $(x,y)$ in the NF.

### v2

@@ I Introduction
The use of entangled photons has also been extended to phase imaging. Phase imaging plays a pivotal role in revealing with enhanced clarity, the subtle variations in the optical path length of light passing through a transparent sample and has found widespread usage in biomedical science Park et al. (2018); Micó et al. (2019); Nguyen et al. (2022). Interferometry based quantum phase retrieval techniques includes the use of photonic N00N states Ono et al. (2013); Israel et al. (2014) and the interference between successive laser passes of spontaneous parametric down-conversion (SPDC) events Black et al. (2023). Phase imaging has also been demonstrated through holography with entangled photons Defienne et al. (2021b); Thekkadath et al. (2023). Non-interference methods includes ghost imaging approaches Abouraddy et al. (2004), the use of transport of intensity equation between intensity images obtained from different image planes Lu et al. (2015); Ortolano et al. (2023) and through the simultaneous measurement of the position and momentum information of photons through a Fourier Ptychography approach Aidukas et al. (2019) and differential phase gradient approach Hodgson et al. (2023). These techniques offers various advantages such as sub-shot-noise sensitivity Ono et al. (2013); Israel et al. (2014); Ortolano et al. (2023); Lu et al. (2015), strong background-noise tolerance Aidukas et al. (2019); Defienne et al. (2021b) and ease of use, such as phase stability Thekkadath et al. (2023) or allowing scanning-free measurement compared to the classical counterpart Hodgson et al. (2023).
@@ II Experimental Setup
The setup of our quantum phase microscope is depicted in Fig. 1. A 20 mW, 405 nm continuous-wave (CW) laser, with a 1 mm beam diameter, is used to pump a 1 mm thick Type II periodically-poled potassium titanyl phosphate (ppKTP) crystal to generate, through the process of spontaneous parametric down-conversion (SPDC), orthogonally polarized photon pairs at 810 nm that are correlated in time and entangled in the position-momentum degrees of freedom. The photon pairs are directed through a 4f-imaging system to illuminate a phase target placed in the NF plane of the nonlinear crystal. The photons are then separated using a polarizing beamsplitter. The NF of the phase target is imaged onto a time-tagging camera (TPX3CAM Nomerotski (2019); ASI ) through the signal photons, while the FF of the target is projected onto a corner of the camera through a separate path using the idler photons. To identify photon pairs, time correlation measurement is conducted with a coincidence window of 20 ns. Because the photons are position-momentum entangled, we can infer the momentum of each signal photon detected at a NF position from the partnered idler photon, whose momentum is measured in the FF. Thus, 5-dimensional information, in time, position and momentum, is obtained for each photon.
@@
Before conducting these measurements with the phase target, a one-time reference measurement without the phase target in place is performed to determine $\phi_{0}(x,y)$. Note that background contributions has not been taken into account in eq. 1, see the Supplementary on the modifications required for background correction.
III Experimental Results
III.1 Quantitative phase measurement
@@ III.2 Background mitigation
Here, we simulate the effect of a dynamic background by introducing numerically generated photon detection events from a Gaussian shaped beam profile into the raw camera data, both in the NF and FF. The properties of this background, including its intensity, width, and position, were all randomized over time. Refer to the Supplementary for more details on the background generation.
@@
However, even after coincidence detection, an uncorrelated background remains due to photons that accidentally arrive within the coincidence window despite not being generated in the same SPDC event. These are due to accidental coincidences between a SPDC photon and a background photon, two background photons, or two SPDC photons belonging to different pair events. In the second stage, the rate of accidental coincidences are measured by selecting a coincidence window that is outside of the true coincidence window, in this case, 50 ns before it. The contribution of background coincidences can then be subtracted from the the centroid shift to mitigate the effect of a dynamic background on the recovered phase profile. Refer to the Supplementary for more details on the detection and subtraction of background light.
@@
III.3 Reduced far-field resolution
@@ IV Discussion
During the submission of this work, we are made aware of a work using a similar phase gradient approach but using a SH sensor instead of a time-tagging camera for the measurement of biphoton spatial wave function Zheng et al. (2024).